### v2

# Multimodal Fact Checking with Unified Visual, Textual, and Contextual Representations

Abstract
The growing rate of multimodal misinformation, where claims are supported by both text and images poses significant challenges to fact-checking systems that rely primarily on textual evidence. In this work, we propose a unified framework for fine-grained multimodal fact verification called MultiCheck, designed to reason over structured textual and visual signals. Our architecture combines dedicated encoders for text and images with a fusion module that captures cross-modal relationships using element-wise interactions. A classification head then predicts the veracity of a claim, supported by a contrastive learning objective that encourages semantic alignment between claim-evidence pairs in a shared latent space. We evaluate our approach on the Factify 2 dataset, achieving a weighted F1 score of 0.84, substantially outperforming the baseline. These results highlight the effectiveness of explicit multimodal reasoning and demonstrate the potential of our approach for scalable and interpretable fact-checking in complex, real-world scenarios.
1 Introduction:
Misinformation has become a serious concern in today’s digital environment, affecting many areas like politics, public health, and finance [1]. While early instances of false information were mostly text-based [2, 3, 4], modern misinformation campaigns increasingly blend text with images, audio, and video making them more persuasive and harder to detect [5, 6, 7]. This rise in multimodal misinformation reveals the limitations of traditional fact-checking systems, which primarily focus on textual content [8, 9, 6]. In response, the research community has turned to multimodal fact-checking, where claims are verified using both textual and visual contents [10, 9]. Recent benchmarks such as FEVER [11], Mocheg [12], and Factify 2 [13] have helped advance research in this direction, see section 2 for further details. Multimodal fact-checking remains a difficult and open challenge. This is because, the task goes beyond simple classification. It requires understanding how modalities interact and whether they support or contradict each other. For example in Figure 1, where a claim is paired with a scientific image that appears credible but is actually unrelated. Identifying this mismatch requires more than surface-level alignment. Thus we need a structured reasoning across both textual and visual modalities.
Figure 1: Refuting a viral claim using combined text and image evidence.
Existing methods relied on one of two strategies: (i) concatenating image-text embeddings  [14, 15], or (ii) joining vision and language encoders through late fusion [16]. For example, Mocheg[12] processes claims and evidence through modality-specific encoders and fuses their outputs without deep cross-modal interaction, whereas Pro-CoFactv2[17] leverages attention-based mechanisms. While these strategies capture basic correlations, they often fail to detect fine contradictions or poorly aligned semantics. we believe they fail to separate modality-specific support signals and make the reasoning process transparent. We, on the other hand, have proposed a unified fact-checking framework called "MultiCheck", where we used (i) a novel fusion technique, which captures semantic relation via element-wise difference and product operations, and (ii) a contrastive learning objective that aligns semantically similar claim-document pairs in a shared latent space, improving representation consistency across modalities. Our approach is inspired by prior advances in multimodal relational reasoning. The element-wise difference and product were shown to be capturing fine-grained interactions between paired inputs in natural language inference models [18] and bilinear attention models [19]. These operations encode both alignment and divergence between two modalities, enabling more expressive cross-modal representations. To further strengthen semantic alignment, we integrated a contrastive head that serves the objective of contrastive learning that operates on projected claim-document embeddings. This module is trained with a symmetric InfoNCE loss [20], encouraging the model to align semantically related pairs while pushing apart unrelated ones. Unlike prior methods that used frozen embeddings or shallow probes [21], our model is fully trainable end-to-end and jointly optimizes for both classification and contrastive learning, resulting in a more discriminative and robust representation for multimodal fact verification.
Our contributions are as follows:
  • We have introduced a unified multimodal fact-checking architecture, "MultiCheck", that combines structured text and image features. It incorporates a contrastive head and a fusion module to align semantically related claim-evidence pairs while separating unrelated ones. As a result, we achieved a new state-of-the-art with the weighted F1 score of 0.84, outperforming the baseline by 27%.
  • We have conducted thorough ablations to assess the impact of different backbones, fusion strategies, and training objectives. The results reveal that relational fusion using difference and product consistently outperforms simple concatenation, and that the inclusion of the contrastive loss significantly boosts performance, particularly in ambiguous or weak evidence scenarios.
  • We have performed a comprehensive error analysis using statistical significance tests, showing that our model not only outperforms the baseline but does so in a structurally meaningful way, correcting more errors than it introduces. Discordant pair comparisons and contingency heatmaps analyses reveals consistent improvements across challenging veracity classes.
  • We further supported our findings through qualitative analysis, highlighting how OCR cues and visual metadata can decisively shift predictions in subtle cases often missed by prior systems.
2 Related Work
Recent research in automated fact-checking has highlighted the growing importance of incorporating multiple modalities, to tackle the diverse and evolving forms of misinformation [22]. Early works, such as FEVER [23] and CLEF2018 [24] are primarily focused on verifying textual claims, laying foundational methods for claim verification based solely on textual evidence. However, later studies found that misinformation exploits images, videos, and audio alongside text to build convincing narratives [25, 26]. These studies have revealed the limitations of purely text-based fact-checking methods and sparked a shift toward multimodal fact-checking. To tackle the limitation of text-based methods, systems were designed to jointly process and reason over diverse types of content. For example, several studies, such as [16] and [27, 28], have explored architectures that integrated textual and visual features. These models employ mechanisms like attention or contrastive learning to enhance detection accuracy. Recent work by [29] shows that a probing classifier combining separate text and image embeddings can outperform intrinsic VLM features on datasets like Factify 2. In addition to these developments, comprehensive surveys, such as [10], offers detailed overview regarding the emerging field of multimodal fact-checking. They highlighted both the technical challenges and the promising research directions ahead. Key challenges include aligning information across modalities, managing incomplete or noisy evidence, and ensuring scalability for practical deployment. Despite significant progress, effectively integrating multimodal information remains an open research problem. This challenge continues to motivate the development of new architectures and learning methods. Robust fact verification in multimodal contexts still requires innovative solutions.
3 Dataset details:
We have used the Factify-2 dataset [30] in our experiments. It is a large-scale multimodal dataset with 50,000 annotated claims spanning across domains such as politics, health, and global affairs. Claims were collected from verified news accounts on Twitter and trusted fact-checking platforms such as PolitiFact [^1] https://www.politifact.com, Alt News [^2] https://www.altnews.in, and BoomLive [^3] https://www.boomlive.in. Each sample includes a claim-evidence pair with separate text, image, and OCR text, as shown in Figure 2. Unlike prior datasets such as Weibo [31], FakeNewsNet [32], and Mocheg [33], Factify-2 provides both text and image modalities at the claim level, enabling fine-grained multimodal verification. Claims are labeled one of the five labels based on their relationship with the evidence. Since the original test set was not publically available, we followed split protocol of prior work [29], by using the original validation set for testing and reallocating 7,500 training samples for validation. Table 1 shows the class distributions.
Figure 2: Example of a sample from the dataset.
Factify 2 [30]
Classes Train Validation Test Total
Support_Multimodal 5,580 1,420 1,500 8,500
Support_Text 5,485 1,515 1,500 8,500
Insufficient_Multimodal 5,472 1,528 1,500 8,500
Insufficient_Text 5,494 1,506 1,500 8,500
Refute 5,469 1,531 1,500 8,500
Total 27,500 7,500 7,500 42,500
Table 1: Dataset statistics for Factify 2 across training, validation, and test splits.
4 Methodology
In this section, we have reported our proposed framework, schematically depicted in Figure 4. Our framework has four components, i.e. (i) text module, (ii) image module, (iii) fusion module, and (iv) classification module. Each component of our framework is described as follows:
Text module: This module converts the claims, associated evidence documents and OCR texts extracted from images into an embedding space. For each instance, we concatenate the claim text with the OCR text from associated images to build a richer, context-aware representation. This fused input helps capture nuances that may not be evident in the claim alone. For example, in Figure 3, the OCR content reveals contradicting cues absent from the main claim, enabling more accurate veracity detection. We have used pre-trained language models such as RoBERTa [34], DeBERTa [35], or SBERT [36] for this purpose. This yields token IDs and attention masks of shape $\mathbb{R}^{b\times L}$, where $b$ is the batch size and $L$ is the sequence length. Each sequence is processed through a shared encoder to produce contextual embeddings:
$\mathbf{H}_{\text{text}}\in\mathbb{R}^{b\times L\times d_{\text{text}}}$
Here, $d_{\text{text}}$ is the hidden size of the language model (e.g., 1024 for RoBERTa and DeBERTa, 768 for SBERT). We extract the [CLS] token embedding as a compact sequence representation:
$\mathbf{h}_{[\text{CLS}]}\in\mathbb{R}^{b\times d_{\text{text}}}$
To align this with other modalities, we have projected the [CLS] embedding into a shared latent space using a linear transformation:
$\mathbf{V}_{\text{text}}=\mathbf{W}_{\text{text}}\cdot\mathbf{h}_{[\text{CLS}]}+\mathbf{B},\quad\mathbf{V}_{\text{text}}\in\mathbb{R}^{b\times h}$
Here, $h$ denotes the latent dimension, $\mathbf{W}_{\text{text}}$ is the projection weight matrix, and $\mathbf{B}$ is a bias term.
This module ensures that semantic and contextual information from both textual and OCR channels is preserved and embedded into a format suitable for cross-modal fusion in downstream stages.
Figure 3: Illustration: The claim suggests Nepal shot down an Indian helicopter. However, the OCR text contradicts this by suggesting Indian aggression, not Nepali. Without OCR, the model could misclassify this. The fused text representation enables correct "Refute" labeling.
Figure 4: Intuitive fusion representation using element-wise difference and product.
Image module: The image module encodes the visual evidence associated with claims and supporting evidence documents into a shared embedding space. Each image is first preprocessed and passed through a shared vision encoder either ResNet50 [37] or Vision Transformer (ViT) [38]. This produces visual features of shape $\mathbf{f}_{\text{img}}\in\mathbb{R}^{b\times d_{\text{img}}}$, where $b$ is the batch size and $d_{\text{img}}$ corresponds to the raw output dimension (2048 for ResNet and 768 for ViT). To align with the text representation space, we project these features into a shared latent space using a linear transformation:
$\mathbf{V}_{\text{img}}=\mathbf{W}_{\text{img}}\cdot\mathbf{f}_{\text{img}}+\mathbf{B},\quad\mathbf{V}_{\text{img}}\in\mathbb{R}^{b\times h}$
Here, $\mathbf{W}_{\text{img}}$ is the projection matrix, $\mathbf{B}$ is the bias term, and $h$ denotes the shared latent dimensionality. To ensure compatibility for multimodal reasoning, both text and image embeddings are projected into the same vector space through modality-specific linear layers. At the end of this module, we obtain four representative vectors of shape $\mathbb{R}^{b\times h}$: $\mathbf{V}_{\text{claim text}}$ and $\mathbf{V}_{\text{doc text}}$ from the text module, and $\mathbf{V}_{\text{claim img}}$ and $\mathbf{V}_{\text{doc img}}$ from the image module. These embeddings serve as the inputs for the fusion and classification modules that follow.
Fusion module: Our key innovation lies in the fusion module. It is designed to capture both direct and relational interactions between multimodal evidence. Prior systems rely on simple feature concatenation [39, 40, 41]. In contrast, our framework explicitly captures alignment and divergence between claim and document evidence. It uses element-wise difference and product operations over their multimodal representations embeddings. This approach is inspired by relational reasoning techniques shown effective in prior research [19, 18, 14, 42, 43].
Specifically, We first generated two integrated multimodal representations by concatenating text and image embeddings for the claim and the document:
$\mathbf{V}_{\text{claim repr}}$ $=\mathbf{V}_{\text{claim text}}\oplus\mathbf{V}_{\text{claim img}}$
$\mathbf{V}_{\text{doc repr}}$ $=\mathbf{V}_{\text{doc text}}\oplus\mathbf{V}_{\text{doc img}}$
To extract finer relationships, we then computed (i) element-wise difference: which highlights differences between claim and document representations, and (ii) element-wise product: which emphasizes regions of strong alignment, between those two vectors denoted by:
$\mathbf{V}_{\text{diff}}$ $=\left|\mathbf{V}_{\text{claim repr}}\ominus\mathbf{V}_{\text{doc repr}}\right|$
$\mathbf{V}_{\text{prod}}$ $=\mathbf{V}_{\text{claim repr}}\otimes\mathbf{V}_{\text{doc repr}}$
where $\ominus$ and $\otimes$ denote element-wise difference and product, respectively. It helps in determining, whether the multimodal input supports, refutes, or fails to verify a claim. Finally, these four vector embeddings: the claim representation, evidence document representation, difference, and product are concatenated, denoted by:
$\mathbf{V}_{\text{fused}}=\mathbf{V}_{\text{diff}}\oplus\mathbf{V}_{\text{prod}}\oplus\mathbf{V}_{\text{claim repr}}\oplus\mathbf{V}_{\text{doc repr}}$
Where $\oplus$ shows the operation of concatenation. The fused vector is passed through a feedforward network with GELU activation [44], and dropout, yielding the final fused vector used for classification, denoted by:
$\mathbf{V}_{\text{final}}=\text{FFN}(\mathbf{V}_{\text{fused}})\in\mathbb{R}^{b\times h}$
Where h denotes the dimensionality of the shared latent space and b shows the batch-size. This final fused representation captures refined multimodal interactions critical for reliable fact verification. To complement the supervised classification objective, we have incorporated a contrastive projection head that encourages semantic alignment between matching claim-document pairs. Each multimodal representation is passed through a shared two-layer projection network:
$\mathbf{Z}_{\text{claim}}$ $=f_{\text{proj}}(\mathbf{V}_{\text{claim repr}})$
$\mathbf{Z}_{\text{doc}}$ $=f_{\text{proj}}(\mathbf{V}_{\text{doc repr}})$
where the projection function $f_{\text{proj}}$ is defined as:
$f_{\text{proj}}(\mathbf{v})=\mathbf{W}_{2}\cdot\text{ReLU}(\mathbf{W}_{1}\cdot\mathbf{v}+\mathbf{B}_{1})+\mathbf{B}_{2}$
where $\mathbf{v}$ is the concatenated multimodal representation of either claim or the evidence document, $\mathbf{W}_{1}\in\mathbb{R}^{h\times 2h}$ and $\mathbf{W}_{2}\in\mathbb{R}^{h\times h}$, and $\mathbf{B}_{1}$, $\mathbf{B}_{2}$ are bias terms. This bottleneck structure reduces the concatenated vector from 2h to h and reprojects it, enabling richer discriminative representations. The resulting embeddings, $\mathbf{Z}_{\text{claim}}$ and $\mathbf{Z}_{\text{doc}}$, are used during training to compute a symmetric InfoNCE loss [20]. This loss encourages the model to pull together representations of matching claim-document pairs while pushing apart those of unrelated ones. By aligning semantically similar inputs across modalities, the contrastive learning objective improves the model’s sensitivity to subtle semantic differences across modalities and strengthens its generalization capabilities.
Classification module: The classification module translates the fused multimodal representation into a final prediction over five fine-grained veracity classes. The final fused vector $\mathbf{V}_{\text{final}}\in\mathbb{R}^{b\times h}$ is passed through a linear classifier:
$\mathbf{P}=\mathbf{W}_{\text{cls}}\cdot\mathbf{V}_{\text{final}}+\mathbf{B},\quad\mathbf{P}\in\mathbb{R}^{b\times 5}$
where $b$ is the batch size and $\mathbf{P}$ contains the predicted logits over the five veracity labels. We use standard cross-entropy loss to supervise this classification:
$\mathcal{L}_{\text{CE}}=-\frac{1}{b}\sum_{i=1}^{b}\log\left(\frac{\exp\left(Z_{i,y_{i}}\right)}{\sum_{j=1}^{5}\exp\left(Z_{i,j}\right)}\right)$
To improve discriminative learning and cross-modal alignment, we have incorporated a symmetric contrastive objective. Projected embeddings of the claim and document representations, denoted as $\mathbf{Z}{\text{claim}}$ and $\mathbf{Z}{\text{doc}}$, are normalized:
$\hat{\mathbf{z}}=\frac{\mathbf{z}}{\lVert\mathbf{z}\rVert_{2}}$
We then compute the pairwise similarity matrix $\mathbf{S}$ as:
$\mathbf{S}=\hat{\mathbf{Z}}_{\text{claim}}\cdot\hat{\mathbf{Z}}_{\text{doc}}^{\top}\in\mathbb{R}^{b\times b}$
The $S_{ij}$ represents the cosine similarity between the $i$-th claim and $j$-th document. Then the InfoNCE loss is computed in both directions:
claim $\rightarrow$ evidence document:
$\text{Loss}_{\text{claim}\rightarrow\text{doc}}=-\frac{1}{b}\sum_{i=1}^{b}\log\frac{\exp\left(S_{ii}/\tau\right)}{\sum_{j=1}^{b}\exp\left(S_{ij}/\tau\right)}$
document evidence $\rightarrow$ claim:
$\text{Loss}_{\text{doc}\rightarrow\text{claim}}=-\frac{1}{b}\sum_{j=1}^{b}\log\frac{\exp\left(S_{jj}/\tau\right)}{\sum_{i=1}^{b}\exp\left(S_{ij}/\tau\right)}$
Then the final contrastive loss is the average of both:
$\mathcal{L}_{\text{contrastive}}=\frac{1}{2}\left(\text{Loss}_{\text{claim}\rightarrow\text{doc}}+\text{Loss}_{\text{doc}\rightarrow\text{claim}}\right)$
The total training objective combines classification and contrastive supervision:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda\cdot\mathcal{L}_{\text{contrastive}}$
where $\lambda$ balances the influence of the contrastive signal. We have set $\lambda=0.1$ in all experiments. This dual-objective training not only improves classification accuracy but also structures the latent space such that semantically aligned claim-document pairs are closer, while unrelated pairs remain well-separated. The contrastive head consistently boosts performance across all evaluated models.
5 Experiments:
In this section, we have presented the experiments to evaluate the effectiveness of our framework. We have assessed how learning strategies, architectures, and fusion methods with element-wise operations affect multimodal performance.
5.1 Experimental setup:
All experiments were conducted on a single NVIDIA A100 80GB GPU using PyTorch and Hugging Face Transformers. We have evaluated model performance using the weighted F1 score, which balances class-wise F1 scores based on support (i.e., true instance count per class). This metric provides a fairer summary under label imbalance compared to unweighted F1. Formally, if $F_{i}$ is the F1 score and $n_{i}$ is the support for class $i$, then:
$\text{Weighted-F}_{1}\;=\;\frac{\sum_{i}n_{i}\,F_{i}}{\sum_{i}n_{i}}\,.$
We have performed thorough hyperparameter tuning for both the contrastive and non-contrastive variants. The final configuration for the experiments are reported in Table2. We have experimented with a range of pre-trained language models for text and backbone vision encoders for image features. Both modalities are projected into a shared latent space. Code and trained models will be released for reproducibility.
Component Hyperparameter Value
Reproducibility Seeds 42, 57, 196, 906
Tokenization Max length 128
Contrastive Loss Temperature ($\tau$) 0.1
Loss weight ($\lambda$) 0.1
Optimization Optimizer Adam
Learning rate $1\times 10^{-5}$
Batch size 32
Num workers 4
Epochs 20
LR Scheduling Scheduler ReduceLROnPlateau
Early Stopping Patience 5
Table 2: Hyperparameter settings used in all experiments.
5.2 Baselines:
For comparative evaluation, we have reproduced the Pro_CoFactv2 model, originally proposed by [45], which achieves state-of-the-art performance on the Factify 2 benchmark. This model leverages large pretrained encoders: DeBERTa-large for text and SwinV2-base for images along with adapter-based projection layers and a lightweight fusion classifier. Both encoders are unfrozen and trained end-to-end. Pro_CoFactv2 applies linear fusion of text and image features, and jointly optimizes a multi-loss objective: cross-entropy for classification and a supervised contrastive loss scaled by 0.3 to improve intra-class alignment in the embedding space. We have replicated the original configuration using the same random seeds (42, 57, 196, 906), learning rate ($5\times 10^{-5}$), batch size (32), and training duration (20 epochs). Sequence length is set to 128 due to GPU constraints, with a dropout rate of 0.1, 12 attention heads, and an intermediate hidden dimension of 256. Experiments were conducted on a single NVIDIA A100 80GB GPU. Reproduced performance is reported in Table 3.
Class Precision Recall F1 Score
Support Text 0.48 (±0.04) 0.38 (±0.06) 0.42 (±0.03)
Support Multimodal 0.50 (±0.04) 0.61 (±0.02) 0.55 (±0.02)
Insufficient Text 0.50 (±0.01) 0.44 (±0.07) 0.46 (±0.04)
Insufficient Multimodal 0.43 (±0.02) 0.46 (±0.06) 0.44 (±0.02)
Refute 0.98 (±0.00) 0.98 (±0.00) 0.98 (±0.00)
Weighted F1 Score 0.57 (±0.01)
Table 3: Performance of the reproduced baseline model (Pro-CoFactv2) on the Factify 2 dataset.
5.3 Experimental Variants:
In addition to evaluating the baseline, we conducted experiments to assess the impact of our approach by comparing two main model variants:
With Contrastive Head: This version integrates a contrastive projection head applied to the multimodal representations of claims and evidence documents. Training includes a contrastive loss in addition to the standard cross-entropy loss. This encourages the model to learn modality-consistent and semantically aligned embeddings.
Without Contrastive Head: In this version, the contrastive projection head and the associated loss are omitted. The model relies solely on cross-entropy loss applied to the fused multimodal representation for classification. Results are presented in Tables 4, 5, with discussion in Section 6.
6 Results and Discussion:
Models w/ Contrastive Head Support Text Support Multimodal Insufficient Text Insufficient Multimodal Refute Weighted F1 Score
Roberta + ResNet50 0.77 (±0.02) 0.83 (±0.01) 0.82 (±0.00) 0.76 (±0.02) 1.00 (±0.00) 0.84 (±0.01)
Roberta + ViT 0.77 (±0.02) 0.84 (±0.01) 0.82 (±0.00) 0.77 (±0.01) 1.00 (±0.00) 0.84 (±0.01)
DeBERTa + ViT 0.77 (±0.01) 0.84 (±0.01) 0.83 (±0.00) 0.78 (±0.01) 1.00 (±0.00) 0.84 (±0.01)
DeBERTa + ResNet50 0.76 (±0.02) 0.81 (±0.02) 0.81 (±0.01) 0.75 (±0.02) 1.00 (±0.00) 0.83 (±0.01)
SBERT + ResNet50 0.75 (±0.01) 0.83 (±0.01) 0.79 (±0.00) 0.76 (±0.01) 1.00 (±0.00) 0.82 (±0.00)
Baseline 0.42 (±0.03) 0.55 (±0.02) 0.46 (±0.04) 0.44 (±0.02) 0.98 (±0.00) 0.57 (±0.01)
Table 4: Class-wise F1 scores and weighted F1 for various model combinations on the Factify 2 dataset with contrastive head. Each value is the mean F1 ±std across seeds.
Models w/o Contrastive Head Support Text Support Multimodal Insufficient Text Insufficient Multimodal Refute Weighted F1 Score
Roberta + ResNet50 0.74 (±0.01) 0.82 (±0.01) 0.78 (±0.01) 0.75 (±0.01) 0.99 (±0.01) 0.82 (±0.01)
Roberta + ViT 0.75 (±0.01) 0.82 (±0.01) 0.79 (±0.01) 0.75 (±0.01) 1.00 (±0.00) 0.82 (±0.00)
DeBERTa + ViT 0.74 (±0.01) 0.82 (±0.00) 0.78 (±0.01) 0.75 (±0.01) 1.00 (±0.00) 0.82 (±0.00)
DeBERTa + ResNet50 0.74 (±0.01) 0.81 (±0.01) 0.78 (±0.01) 0.74 (±0.01) 1.00 (±0.00) 0.81 (±0.00)
SBERT + ResNet50 0.71 (±0.03) 0.81 (±0.02) 0.75 (±0.02) 0.74 (±0.02) 0.99 (±0.00) 0.78 (±0.05)
Baseline 0.42 (±0.03) 0.55 (±0.02) 0.46 (±0.04) 0.44 (±0.02) 0.98 (±0.00) 0.57 (±0.01)
[…]
We have evaluated our proposed framework on Factify 2 using various language and vision encoder backbones. Performance is measured via weighted F1 across five veracity labels. As shown in Tables 4 and 5, all variants significantly outperforms the baseline, models integrated with contrastive-head consistently outperform their non-contrastive counterparts. Specifically gains are most notable in “Insufficient” and “Refute” cases. This highlights the effectiveness of our fusion and contrastive learning objective in aligning text and image evidence.
6.1 Insights:
Contrastive learning boosts accuracy: Across all settings, the inclusion of contrastive supervision yields statistically significant gains. It consistently outperforms non-contrastive counterparts, see Table 6. The consistent asymmetry in discordant prediction counts, as shown in Figure 5, confirms that contrastive models correct significantly more baseline errors than they introduce.
Role of fusion strategy: The fusion module using element-wise difference and product provides critical advantages over simple concatenation. These operations help the model clearly match or contrast the claim with the evidence, improving its reasoning.
Model robust nature: Whether paired with ResNet50 or ViT, and regardless of the language encoder, the contrastive head along with fusion module consistently improves performance, highlighting its modularity and general applicability.
7 Error analysis:
Models w/ Contrastive Head
McNemar’s
$\chi^{2}$
McNemar’s
p-value
Significance at $\alpha=0.05$
Bowker’s
$\chi^{2}$ (df=10)
Bowker’s
p-value
Reject
symmetry
Roberta + ResNet50 1238.56 $\ll 0.05$ 91.85 $2.33\times 10^{-15}$
Roberta + ViT 1333.14 $\ll 0.05$ 132.10 $0.00$
DeBERTa + ViT 1480.57 $\ll 0.05$ 118.33 $0.00$
DeBERTa + ResNet50 1217.24 $\ll 0.05$ 166.16 $0.00$
SBERT + ResNet50 1070.45 $\ll 0.05$ 82.030 $2.00\times 10^{-13}$
Table 6: Significance-test results comparing the multimodal model against the baseline on Factify 2.
Models w/o Contrastive Head
McNemar’s
$\chi^{2}$
McNemar’s
p-value
Significance at $\alpha=0.05$
Bowker’s
$\chi^{2}$ (df=10)
Bowker’s
p-value
Reject
symmetry
Roberta + ResNet50 1129.91 $\ll 0.05$ 75.64 $3.56\times 10^{-12}$
Roberta + ViT 1243.45 $\ll 0.05$ 140.44 $0.00$
DeBERTa + ViT 965.16 $\ll 0.05$ 122.93 $0.00$
DeBERTa + ResNet50 1195.33 $\ll 0.05$ 153.76 $0.00$
SBERT + ResNet50 891.88 $\ll 0.05$ 100.06 $0.00$
Table 7: Significance-test results comparing the multimodal model against the baseline on Factify 2.
To better understand the behavioral differences between our models and the baseline, we performed a detailed error analysis using both statistical significance tests and qualitative assessments.
Statistical comparisons: we have applied McNemar’s test for overall accuracy and Bowker’s test for class-level shifts. Tables 67 show consistently high $\chi^{2}$ scores with p-values < 0.05, confirming significant improvements. Whereas Bowker’s results further indicate structured, non-random gains in class predictions.
Discordant pair analysis: We compared our model predictions over the baseline using McNemar’s 2×2 contingency setup to identify disagreement cases. Here, (b) counts instances where the baseline is correct and our model is errors, while (c) captures the opposite. In all five model variants, (c) consistently exceeds (b), as shown in Figure 5, indicating that our model reliably corrects baseline errors. These findings align with significance tests, further validating the effectiveness of our approach.
Qualitative Analysis: To better understand model behavior, we analyzed the mismatches between our approach and the baseline. As shown in Figure 6,the baseline relied only on text, predicting Insufficient Text. However, Our model used the OCR credit “Helen Sloan/HBO” to recognize the image as a licensed promotional photograph, detail suggesting the image does not contribute new factual content, and correctly predicting Insufficient Multimodal. By recognizing the lack of substantive support in both text and image, our model rightly predicted Insufficient Multimodal. This highlights the role of OCR-extracted image sources in improving factual reasoning. Additional examples are discussed below:
Additional examples of qualitative analysis:
Here, we have presented the dataset-referenced examples comparing our model with the baseline, showing how OCR cues such as credits, agency marks, or overlays can clarify or complicate the multimodal verification.
Example A: When MultiCheck outperforms the baseline
  • ID 4681: The baseline, relying only on text, misclassified “Made in India vaccines” as Support_Text. In contrast, our model’s multimodal fusion leveraged the OCR cue “COLO STORAE” (cold storage) to detect ambiguity and correctly predicted Insufficient_Text. This illustrates how even noisy OCR can reveal context missed by text-only models.
  • ID 6968: The baseline predicted Support_Text based on name and context matches in text (e.g., “Governor Jagdeep Dhankar”). In contrast, our model used the OCR cue “ANI” to identify the image as a generic press photo and, combining it with weak textual alignment, correctly predicted Insufficient_Multimodal. This demonstrates its ability to dismiss superficial visual signals.
Example B: When baseline outperforms MultiCheck
  • ID 7171: In this case, our model was overly cautious. The OCR tag “ANI” led it to dismiss the image as generic, resulting in an Insufficient_Multimodal prediction. However, the textual portion “will lay foundation stone” is a direct and verifiable event announcement, and the image provided valid context. The baseline correctly predicted Support_Multimodal, highlighting that OCR cues can sometimes mislead when visual content, though generic, remains contextually relevant.
    (a) Roberta+ResNet50 (with contrastive head)
    (b) Roberta+ResNet50 (with contrastive head)
    (c) Roberta+ResNet50 (without contrastive head)
    (d) Roberta+ResNet50 (without contrastive head)
    (e) Roberta + ViT (with contrastive head)
    (f) Roberta + ViT (with contrastive head)
    (g) Roberta + ViT (without contrastive head)
    (h) Roberta + ViT (without contrastive head)
    (i) DeBERTa + ViT (with contrastive head)
    (j) DeBERTa + ViT (with contrastive head)
    (k) DeBERTa + ViT (without contrastive head)
    (l) DeBERTa + ViT (without contrastive head)
    (m) DeBERTa + ResNet50 (with contrastive head)
    (n) DeBERTa + ResNet50 (with contrastive head)
    (o) DeBERTa + ResNet50 (without contrastive head)
    (p) DeBERTa + ResNet50 (without contrastive head)
    (q) SBERT + ResNet50 (with contrastive head)
    (r) SBERT + ResNet50 (with contrastive head)
    (s) SBERT + ResNet50 (without contrastive head)
    (t) SBERT + ResNet50 (without contrastive head)
    Figure 5: Comparison of discordant-pair barplots (first column per model) and McNemar 2×2 contingency heatmaps (second column per model), both with and without the contrastive head. Models (top to bottom): Roberta + ResNet50, Roberta + ViT, DeBERTa + ViT, DeBERTa + ResNet50, SBERT + ResNet50.
Key Observations:
  • OCR text can be highly informative especially when images include meta-tags, banners, or visual overlays not present in article text.
  • Failure cases arise when OCR includes irrelevant or misleading tokens, causing the model to over- or under-attend to visuals.
  • Future directions: We plan to incorporate an OCR quality gating mechanism and synthetic noisy-OCR augmentation to improve model robustness.
Figure 6: Example of qualitative analysis, Sample from the dataset.
8 Conclusion:
This paper presents a unified framework for fine-grained multimodal fact-checking that jointly reasons over text and images. It integrates structured representations from pre-trained encoders through a relational fusion module and enhances cross-modal alignment via contrastive learning. This enables the model to capture subtle agreements and contradictions between claims and evidence. Our approach consistently outperforms the baseline, with significant gains in complex classes like "Insufficient" and "Refute". Statistical and qualitative analyses confirms that these improvements are consistent and non-random. Overall, these results highlights the value of explicit cross-modal alignment and representation learning for reliable fact verification.
9 Limitation:
While our approach shows strong empirical performance, several limitations remain:
  • Dependence on OCR quality: The model relies on OCR-extracted text, which can vary in quality. Noisy or misleading OCR may cause incorrect predictions due to false visual-textual alignment.
  • No evidence retrieval component: Our framework assumes relevant text and image evidence is already provided. It does not include a retrieval mechanism to source evidence from external knowledge bases or the web.
  • Limited visual understanding: Although pre-trained vision encoders are used, the model lacks deeper visual reasoning, such as object detection, scene understanding, or temporal interpretation.
  • Restricted modalities: The system currently supports only text and image. It does not handle audio, video, or temporal misinformation, which are increasingly common in real-world settings.